# Towards Domain Adaptive Neural Contextual Bandits

### v2

@@ Abstract
Contextual bandit algorithms are essential for solving real-world decision making problems. In practice, collecting a contextual bandit’s feedback from different domains may involve different costs. For example, measuring drug reaction from mice (as a source domain) and humans (as a target domain). Unfortunately, adapting a contextual bandit algorithm from a source domain to a target domain with distribution shift still remains a major challenge and largely unexplored. In this paper, we introduce the first general domain adaptation method for contextual bandits. Our approach learns a bandit model for the target domain by collecting feedback from the source domain. Our theoretical analysis shows that our algorithm maintains a sub-linear regret bound even adapting across domains. Empirical results show that our approach outperforms the state-of-the-art contextual bandit algorithms on real-world datasets.
@@ 2 Related Work
Contextual Bandits. In the realm of adaptive decision-making, contextual bandit algorithms, epitomized by LinUCB (Li et al., 2010), have carved a niche in efficiently balancing the exploitation-exploration paradigm across a spectrum of use cases, notably in complex adaptive systems such as recommender systems (Li et al., 2010, Wang et al., 2022). These algorithmic frameworks, along with their myriad adaptations (Yue & Guestrin, 2011, Agarwal et al., 2014, Li et al., 2016, Korda et al., 2016, Kveton et al., 2017, Foster et al., 2018, Zhou et al., 2019, Mahadik et al., 2020, Xu et al., 2020), have decisively outperformed conventional bandit models that operate devoid of contextual awareness (Auer, 2002). This superiority is underpinned by theoretical models, paralleling the insights in (Auer, 2002), where LinUCB variants are validated to conform to optimal regret boundaries in targeted scenarios (Chu et al., 2011). However, a discernible gap in these methodologies is their reduced efficacy in a new domain, particularly when getting feedback involves high cost. Our proposed DABand bridges this gap by adeptly aligning both source-domain and target-domain representations in the latent space, thereby reducing performance drop when transfer across different domains.
@@
Domain Adaptation. The landscape of domain adaptation has been extensively explored, as evidenced by a breadth of research (Pan & Yang, 2009, Pan et al., 2010, Long et al., 2018, Saito et al., 2018, Sankaranarayanan et al., 2018, Zhang et al., 2019, Peng et al., 2019, Chen et al., 2019, Dai et al., 2019, Wang et al., 2020, Nguyen-Meidine et al., 2021, Xu et al., 2023). The primary objective of these studies has been the alignment of source and target domain distributions to facilitate the effective generalization of models trained on labeled source data to unlabeled target data. This alignment is conventionally attained either through the direct matching of distributional statistics (Pan et al., 2010, Tzeng et al., 2014, Sun & Saenko, 2016, Peng et al., 2019, Nguyen-Meidine et al., 2021) or via the integration of an adversarial loss (Ganin et al., 2016b, Zhao et al., 2017, Tzeng et al., 2017, Zhang et al., 2019, Kuroki et al., 2019, Chen et al., 2019, Dai et al., 2019, Wang et al., 2020). The latter, known as adversarial domain adaptation, has surged in prominence, bolstered by its theoretical foundation (Goodfellow et al., 2014, Zhao et al., 2018, Zhang et al., 2019, Zhao et al., 2019), the adaptability of end-to-end training in neural network architectures, and its empirical efficacy. However, these methods only work in offline settings, and assumes complete observability of labels in the source domain. Therefore they are not applicable to our online bandit settings.
@@ 3.1 Preliminaries
LinUCB. LinUCB Li et al. (2010) is a classic method for the typical contextual bandit setting. It assumes that the reward is linear to its input context, i.e., ${\bf x}_{i,a}$. During each round $i$, the agent selects an arm according to historical contexts and their corresponding rewards, $r_{i,a}$:
@@
LinUCB with Representation Learning. In this paper, we use Neural LinUCB Xu et al. (2020), Wang et al. (2022) as a backbone model to handle high-dimensional context with representation learning. We assume that there exist a ground-truth encoder $\phi^{*}$ to encode the raw context ${\bf x}_{i,a}$ into a latent-space representation (encoding) $\phi^{*}({\bf x}_{i,a})$. Subsequently, the reward for the context ${\bf x}_{i,a}$ is then $r({\bf x}_{i,a})=\langle\theta^{*},\phi^{*}({\bf x}_{i,a})\rangle$ plus some stochastic noise $\epsilon$. Furthermore, we use the same setting as in Xu et al. (2020) where the ground-truth rewards are restricted to the range of $[0,1]$. Additionally, we further bound predicted rewards by normalizing $\|\widehat{\theta}\|=1$ and $\|\widehat{\phi}({\bf x}_{i,a})\|=1$.
@@ 3.5 Cross-Domain Error Bound for Regression
Prior to introducing our final regret bound, it is necessary to define an additional component. A fundamental challenge in general domain adaptation problems is to manage the divergence between source and target domains. Unfortunately, previous domain adaptation theory only covers classification tasks Ben-David et al. (2010), Zhang et al. (2019). In contrast, the problem of contextual bandits is essentially a reward regression problem. To address this challenge, we introduce the following new definition to bound the error between the source and target domains.
@@
$g\in\mathcal{H}\Delta\mathcal{H}\quad\iff\quad g(x)=|h({\bf x})-h^{\prime}({\bf x})|.$
@@
$\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})$ $=2\enspace\sup_{h,h^{\prime}\in\mathcal{H}}|\mathbb{E}_{x\sim\mathcal{S}}[|h(x)-h^{\prime}(x)|]-\mathbb{E}_{x\sim\mathcal{T}}[|h(x)-h^{\prime}(x)|]|$
@@ 3.6 Final Regret Bound
$\quad+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)}_{Predicted~{}Rewards}},$ (5)
@@
    Typical generalization bounds in DA naively minimize the source regret $R_{{\mathcal{S}}}$ while aligning source and target data in the latent space (i.e., minimizing $\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}$). This is not sufficient, as we need two additional terms, i.e., the regression error $\epsilon_{\mathcal{S}}(h)$ and the predicted reward $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)$, as regularization terms.
@@
    Minimizing the predicted reward $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)$ regularizes the model to avoid overestimating rewards.
@@
  • Predicted Reward. $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)$ serves as a regularization term to regularize the model to avoid overestimating rewards.
The results of the ablation study in Table 4 in Sec. 5 highlight the significance of these two terms. Details for more analysis are in Appendix G.4.1.
@@ 4 Method
Source Regret. Inspired by Neural-LinUCB (Xu et al., 2020), we use LinUCB Li et al. (2010) to update $\widehat{\theta}$ and the following empirical loss function when updating the encoder parameters $\widehat{\phi}$ by back-propagation in round $i$:
@@
13:           Update encoder $\widehat{\phi}_{i}$ and discriminator $g$ by back-propagation to solve the minimax optimization in Eqn. (9).
@@
$\mathcal{L}^{div}=\max_{g}N\cdot(\mathbb{E}^{\mathcal{S}}[\mathcal{L}_{D}(g(\widehat{\phi}({\bf x})),0)]+\mathbb{E}^{\mathcal{T}}[\mathcal{L}_{D}(g(\widehat{\phi}({\bf x})),1)]),$
@@
Predicted Reward. The predicted reward term in Eqn. (5), i.e., $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)$, can be translated to the loss term $\sum\nolimits_{i=1}^{N}\mathcal{L}_{i}^{reg}$, where for each round $i$ we have
@@ 6 Conclusion
In this paper, we introduce a novel domain adaptive neural contextual bandit algorithm, DABand, which adeptly combines effective exploration with representation alignment, utilizing unlabeled data from both source and target domains. Our theoretical analysis demonstrates that DABand can attain a sub-linear regret bound within the target domain. This marks the first regret analysis for domain adaptation in contextual bandit problems incorporating deep representation, shallow exploration, and adversarial alignment. We show that all these elements are instrumental in the domain adaptive bandit setting on real-world datasets. Moving forward, interesting future research could be to uncover more innovative techniques for aligning the source and target domains (mentioned in Appendix F), particularly within the constraints of: (1) bandit settings, as opposed to classification settings, and (2) the high-dimensional and dense nature of the target domain, contrasted with the sparse and simplistic nature of the source domain.

### v3

@@ Abstract
Contextual bandit algorithms are essential for solving real-world decision making problems. In practice, collecting a contextual bandit’s feedback from different domains may involve different costs. For example, measuring drug reaction from mice (as a source domain) and humans (as a target domain). Unfortunately, adapting a contextual bandit algorithm from a source domain to a target domain with distribution shift still remains a major challenge and largely unexplored. In this paper, we introduce the first general domain adaptation method for contextual bandits. Our approach learns a bandit model for the target domain by collecting feedback from the source domain. Our theoretical analysis shows that our algorithm maintains a sub-linear regret bound even adapting across domains. Empirical results show that our approach outperforms the state-of-the-art contextual bandit algorithms on real-world datasets. Code will soon be available at https://github.com/Wang-ML-Lab/DABand.
@@ 2 Related Work
Contextual Bandits. In the realm of adaptive decision-making, contextual bandit algorithms, epitomized by LinUCB (Li et al., 2010), have carved a niche in efficiently balancing the exploitation-exploration paradigm leveraging estimation of reward and uncertainty (Wang et al., 2022; 2016); they have been influential across a spectrum of use cases, notably in complex adaptive systems such as recommender systems (Li et al., 2010, Wang et al., 2022). These algorithmic frameworks, along with their myriad adaptations (Yue & Guestrin, 2011, Agarwal et al., 2014, Li et al., 2016, Korda et al., 2016, Kveton et al., 2017, Foster et al., 2018, Zhou et al., 2019, Mahadik et al., 2020, Xu et al., 2020), have decisively outperformed conventional bandit models that operate devoid of contextual awareness (Auer, 2002). This superiority is underpinned by theoretical models, paralleling the insights in (Auer, 2002), where LinUCB variants are validated to conform to optimal regret boundaries in targeted scenarios (Chu et al., 2011). However, a discernible gap in these methodologies is their reduced efficacy in a new domain, particularly when getting feedback involves high cost. Our proposed DABand bridges this gap by adeptly aligning both source-domain and target-domain representations in the latent space, thereby reducing performance drop when transfer across different domains.
@@
Domain Adaptation. The landscape of domain adaptation has been extensively explored, as evidenced by a breadth of research (Pan & Yang, 2009, Pan et al., 2010, Long et al., 2018, Saito et al., 2018, Sankaranarayanan et al., 2018, Zhang et al., 2019, Peng et al., 2019, Chen et al., 2019, Dai et al., 2019, Wang et al., 2020, Nguyen-Meidine et al., 2021, Xu et al., 2023). The primary objective of these studies has been the alignment of source and target domain distributions to facilitate the effective generalization of models trained on labeled source data to unlabeled target data. This alignment is conventionally attained either through the direct matching of distributional statistics (Pan et al., 2010, Tzeng et al., 2014, Sun & Saenko, 2016, Peng et al., 2019, Nguyen-Meidine et al., 2021) or via the integration of an adversarial loss (Ganin et al., 2016b, Zhao et al., 2017, Tzeng et al., 2017, Zhang et al., 2019, Kuroki et al., 2019, Chen et al., 2019, Dai et al., 2019, Wang et al., 2020, Xu et al., 2022, Liu et al., 2023, Shi & Wang, 2023). The latter, known as adversarial domain adaptation, has surged in prominence, bolstered by its theoretical foundation (Goodfellow et al., 2014, Zhao et al., 2018, Zhang et al., 2019, Zhao et al., 2019), the adaptability of end-to-end training in neural network architectures, and its empirical efficacy. However, these methods only work in offline settings, and assumes complete observability of labels in the source domain. Therefore they are not applicable to our online bandit settings.
Distribution Shift in Bandits. A related study addresses the challenge of policy learning in contextual bandits using historical observational data, particularly when there is a shift in the environment (Si et al., 2023). The authors propose a distributionally robust approach to policy evaluation and learning, ensuring that the learned policy remains effective even under worst-case environmental shifts. However, their focus is distinct from ours. Their goal is to ensure policy robustness against environmental shifts between training and deployment within the same domain, whereas our proposed DABand focuses on adapting policies from a source domain to a different target domain, particularly when direct feedback from the target domain is limited or costly.
@@ 3.1 Preliminaries
LinUCB. LinUCB (Li et al., 2010) is a classic method for the typical contextual bandit setting. It assumes that the reward is linear to its input context, i.e., ${\bf x}_{i,a}$. During each round $i$, the agent selects an arm according to historical contexts and their corresponding rewards, $r_{i,a}$:
@@
LinUCB with Representation Learning. In this paper, we use Neural LinUCB (Xu et al., 2020, Wang et al., 2022) as a backbone model to handle high-dimensional context with representation learning. We assume that there exist a ground-truth encoder $\phi^{*}$ to encode the raw context ${\bf x}_{i,a}$ into a latent-space representation (encoding) $\phi^{*}({\bf x}_{i,a})$. Subsequently, the reward for the context ${\bf x}_{i,a}$ is then $r({\bf x}_{i,a})=\langle\theta^{*},\phi^{*}({\bf x}_{i,a})\rangle$ plus some stochastic noise $\epsilon$. Furthermore, we use the same setting as in Xu et al. (2020) where the ground-truth rewards are restricted to the range of $[0,1]$. Additionally, we further bound predicted rewards by normalizing $\|\widehat{\theta}\|=1$ and $\|\widehat{\phi}({\bf x}_{i,a})\|=1$.
@@ 3.5 Cross-Domain Error Bound for Regression
Prior to introducing our final regret bound, it is necessary to define an additional component. A fundamental challenge in general domain adaptation problems is to manage the divergence between source and target domains. Unfortunately, previous domain adaptation theory only covers classification tasks (Ben-David et al., 2010, Zhang et al., 2019). In contrast, the problem of contextual bandits is essentially a reward regression problem. To address this challenge, we introduce the following new definition to bound the error between the source and target domains.
@@
$g\in\mathcal{H}\Delta\mathcal{H}\quad\iff\quad g({{\bf x}})=|h({\bf x})-h^{\prime}({\bf x})|.$
@@
$\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})$ $=2\enspace\sup_{h,h^{\prime}\in\mathcal{H}}|\mathbb{E}_{{{\bf x}}\sim\mathcal{S}}[|h({{\bf x}})-h^{\prime}({{\bf x}})|]-\mathbb{E}_{x\sim\mathcal{T}}[|h({{\bf x}})-h^{\prime}({{\bf x}})|]|$
@@ 3.6 Final Regret Bound
$\quad+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\sum^{N}_{i=1}\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)}_{Predicted~{}Rewards}},$ (5)
@@
    Typical generalization bounds in DA naively minimize the source regret $R_{{\mathcal{S}}}$ while aligning source and target data in the latent space (i.e., minimizing $\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}$). This is not sufficient, as we need two additional terms, i.e., the regression error $\epsilon_{\mathcal{S}}(h)$ and the predicted reward $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\sum^{N}_{i=1}\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)$, as regularization terms.
@@
    Minimizing the predicted reward $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\sum^{N}_{i=1}\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)$ regularizes the model to avoid overestimating rewards.
@@
  • Predicted Reward. $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\sum^{N}_{i=1}\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)$ serves as a regularization term to regularize the model to avoid overestimating rewards.
The results of the ablation study in Table 4 in Sec. 5 highlight the significance of these two terms. See Appendix G.4 for more discussion on novelty as well as key differences between our DABand and classic domain adaptation.
@@ 4 Method
Source Regret. Inspired by Neural-LinUCB (Xu et al., 2020), we use LinUCB (Li et al., 2010) to update $\widehat{\theta}$ and the following empirical loss function when updating the encoder parameters $\widehat{\phi}$ by back-propagation in round $i$:
@@
13:           Use the Adam optimizer (Diederik, 2014) to update encoder $\widehat{\phi}_{i}$ and discriminator $g$ by back-propagation to solve the minimax optimization in Eqn. (10).
@@
$\mathcal{L}^{div}=\max_{g}N\cdot(\mathbb{E}^{\mathcal{S}}[\mathcal{L}_{D}(g(\widehat{\phi}({\bf x})),0)]+\mathbb{E}^{\mathcal{T}}[\mathcal{L}_{D}(g(\widehat{\phi}({\bf x})),1)]),$ (8)
@@
Predicted Reward. The predicted reward term in Eqn. (5), i.e., $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\sum^{N}_{i=1}\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)$, can be translated to the loss term $\sum\nolimits_{i=1}^{N}\mathcal{L}_{i}^{reg}$, where for each round $i$ we have
@@ 5 Experiments
Performance Analysis of LinUCB and NLinUCB. Our target regret bound in Eqn. (5) consists of 5 terms, i.e., source regret, regression error, data divergence, constant, and predicted reward. After applying linear contextual bandits on the source domain, the source regret enjoys a sub-linear rate. However, in the linear model, the encoder $\widehat{\phi}({\bf x})={\bf x}$ is an identity function and therefore fail to align source-domain and target-domain contexts, leading to an unbounded, large data divergence term in Eqn. (5). This is the main reason why linear contextual bandits, such as LinUCB, perform poorly in the cross-domain contextual bandit setting.
Note that even contextual bandits using a nonlinear encoder, e.g., Neural-LinUCB (NLinUCB), fail in this case because their encoders $\widehat{\phi}({\bf x})$ are not learned to align source-domain and target-domain contexts; they therefore will still lead to an unbounded, large data divergence term in Eqn. (5), and subsequently poor performance in the cross-domain contextual bandit setting.
@@ 6 Conclusion
In this paper, we introduce a novel domain adaptive neural contextual bandit algorithm, DABand, which adeptly combines effective exploration with representation alignment, utilizing unlabeled data from both source and target domains. Our theoretical analysis demonstrates that DABand can attain a sub-linear regret bound within the target domain. This marks the first regret analysis for domain adaptation in contextual bandit problems incorporating deep representation, shallow exploration, and adversarial alignment. We show that all these elements are instrumental in the domain adaptive bandit setting on real-world datasets. Moving forward, interesting future research could be to uncover more innovative techniques for aligning the source and target domains (mentioned in Appendix F), particularly within the constraints of: (1) bandit settings, as opposed to classification settings, and (2) the high-dimensional and dense nature of the target domain, contrasted with the sparse and simplistic nature of the source domain. It would also be interesting to explore the effect of improved uncertainty quantification (Wang & Wang, 2023), potentially via Bayesian deep learning (Wang & Yeung, 2016; 2020, Huang et al., 2019), on DABand’s performance.
Acknowledgments
We sincerely appreciate the generous support from the reviewers/AC for the constructive comments to improve the paper. ZW and XH are partially sponsored by a subcontract of NSF grant 2229876, the A. Russell Chandler III Professorship at Georgia Institute of Technology, and NIH-sponsored Georgia Clinical & Translational Science Alliance. HW is partially supported by the Microsoft Research AI & Society Fellowship, NSF Grant IIS-2127918, NSF CAREER Award IIS-2340125, NIH Grant 1R01CA297832, and the Amazon Faculty Research Award. This research is also supported by NSF National Artificial Intelligence Research Resource (NAIRR) Pilot and the Frontera supercomputer, funded by the National Science Foundation (award NSF-OAC 1818253) and hosted at the Texas Advanced Computing Center (TACC) at The University of Texas at Austin. Finally, we extend our gratitude to the Center for AI Safety (CAIS) for providing the essential computing resources that made this work possible.